-- -- UNIXPROGRAMMER’SMANUAL 4.2BerkeleySoftwareDistribution,Volume2c VirtualVAX−11Version August,1983 ComputerSystemsResearchGroup DepartmentofElectricalEngineeringandComputerScience UniversityofCalifornia Berkeley,California94720 -- -2-Copyright1979,1980RegentsoftheUniversityofCalifornia. Permissiontocopythesedocumentsoranyportionthereofas necessaryforlicenseduseofthesoftwareisgrantedtolicenseesof thissoftware,providedthiscopyrightnoticeandstatementof permissionareincluded. Thedocument‘‘WritingTools−TheSTYLEandDICTION Programs’’iscopyrighted1979byBellTelephoneLaboratories. HoldersofaUNIXTM/32Vsoftwarelicensearepermittedtocopy thisdocument,oranyportionofit,asnecessaryforlicenseduse ofthesoftware,providedthiscopyrightnoticeandstatementof permissionareincluded. Thedocument‘‘TheProgrammingLanguageEFL’’iscopyrighted 1979byBellTelephoneLaboratories.EFLhasbeenapprovedfor generalrelease,sothatonemaycopyitsubjectonlytothe restrictionofgivingproperacknowledgementtoBellTelephone Laboratories. ThismanualreflectssystemenhancementsmadeatBerkeleyand sponsoredinpartbyNSFGrantsMCS-7807291,MCS-8005144, andMCS-74-07644-A04;DOEContractDE-AT03-76SF00034and ProjectAgreementDE-AS03-79ER10358;andbyDefense AdvancedResearchProjectsAgency(DoD)ARPAOrderNo. 4031,MonitoredbyNavalElectronicsSystemsCommandunder ContractNo.N00039-80-K-0649.

-3-UNIXProgrammer’sManual Volume2c−SupplementaryDocuments 4.2BerkeleySoftwareDistribution,VirtualVAX−11Version August,1983 ThisvolumecontainsdocumentswhichsupplementtheinformationinVolume1ofThe UNIX◯rProgrammer’sManual,fortheVirtualVAX-11versionofthesystemasdistributedbyU.C. Berkeley.ThisvolumeisalogicalextensionofVolumes2aand2basprovidedbyBellLaboratories. GeneralWorks 39.BugFixesandchangesin4.2BSD. Abriefdiscussionofthemajoruser-visiblechangesmadetothesystemsincethelast release. GettingStarted 40.AnintroductiontotheCshell Introducingapopularcommandinterpreterandmanyofthecommonlyusedcommands,assuminglittlepriorknowledgeofUNIX. 41.AnIntroductiontoDisplayEditingwithVi Thedocumenttoreadtolearntousetheviscreeneditor. 42.Edit:Atutorial(Revised) Forthosewhopreferlineorientedediting,anintroductionassumingnoprevious knowledgeofUNIXoroftextediting. 43.ExReferenceManual(Version3.1−Oct.1980) Thefinalreferencefortheexeditor,whichunderliesbotheditandvi. 44.ExChanges−Version3.1to3.5 Aquickguidetowhatisnewinversion3.5ofexandvi,forthosewhohaveusedversion3.1. 45.MailReferenceManual(Revised) Completedetailsonthemailprocessingprogram. 46.AGuidetotheDungeonsofDoom(Revised) Anintroductiontothepopulargameofrogue. Languages 47.TheFRANZLISPManual AdialectofLISP,largelycompatiblewithMACLISP. 48.BerkeleyPascalUser’sManual Aninterpretiveimplementationofthereferencelanguage. 49.TheProgrammingLanguageEFL AnintroductiontoapowerfulFORTRANpreprocessorprovidingaccesstoalanguage withstructuresmuchlikeC.

-4-50.BerkeleyFPUser’sManual AdescriptionoftheBerkeleyimplementationofBackus’FunctionalProgramming Language,FP. 51.APortableFortran77Compiler ArevisedversionofthedocumentwhichoriginallyappearedinVolume2b;thisversionreflectstheongoingworkatBerkeley. 52.Introductiontothef77I/OLibrary Adescriptionoftherevisedinput/outputlibraryforFortran77.Thisdocument, whichoriginallyappearedinVolume2b,reflectstheworkcarriedoutatBerkeley. Documentpreparation 53.WritingPaperswithnroffusing−me Apopularmacropackagefornroff. 54.−meReferenceManual Thefinalwordon−me. 55.TheBerkeleyFontCatalog Samplesoffontscurrentlyavailablefortherasterplotters. 56.Writingtools−theStyleandDictionPrograms Descriptionofprogramswhichhelpyouunderstandandimproveyourwritingstyle. 57.Refer−ABibliographySystem Anintroductiontothetoolsusedtomaintainbibliographicdatabases.Themajor program,refer,isusedtoautomaticallyretrieveandformatreferencesbasedondocumentcitations. 58.ARevisedVersionof−ms Aquickdescriptionoftherevisionsmadetothe−msformattingmacrosfornroffand troff. Programming 59.AssemblerReferenceManual Forcompilerwriters. 60.ScreenUpdatingandCursorMovementOptimization Anaideforwritingscreen-oriented,terminalindependantprograms. 61.AnIntroductiontotheSourceCodeControlSystem AusefulintroductoryarticleforthoseuserswhoarelicensedforSCCS. SystemInstallationandAdministration 62.InstallingandOperating4.2BSDontheVAX Thedefinitivereferencedocumentforthoseoccasionswhenyoufindyouneedtostart overagain. 63.Building4.2BSDUNIXSystemswithConfig Anin-depthdiscussionoftheuseandoperationoftheconfigprogram.Thisdocument discusseshowtoconfigureandbuildbinaryimagesofUNIXforyoursite. 64.DiscQuotasinaUNIXEnvironment Alightintroductiontothecareandfeedingofthefacilitieswhichcanbeusedinlimitingdiscresources. 65.4.2BSDLinePrinterSpoolerManual Thisdocumentdescribesthestructureandinstallationprocedureforthelineprinter spoolingsystem. 66.Fsck−TheUNIXFileSystemCheckProgram Areferencedocumentforusewiththefsckprogramduringtimesoffilesystem -5-distress. 67.SendmailInstallationandOperationGuide Thelastwordininstallingandoperatingthesendmailprogram. SupportingDocumentation 68.4.2BSDSystemManual Aconcise,thoughterse,descriptionofthesystemcallinterfaceprovidedin4.2BSD. Thiswillneverbeabestseller. 69.AFastFileSystemforUNIX Adescriptionofthenewfilesystemorganizationdesignandimplementation. 70.4.2BSDNetworkingImplementationNotes Aconcisedescriptionofthesysteminterfacesusedwithinthenetworkingsubsystem. 71.Sendmail−AnInternetworkMailRouter Anoverviewdocumentonthedesignandimplementationofsendmail. 72.MailSystemsandAddressingin4.2BSD Ageneralintroductiontothemailfacilitiesin4.2BSD.